OpenStudy (anonymous):

How do you factor 8z^3 + 1?

OpenStudy (serenity74):

Actually do you know the formula? For the cubic factoring? Cause if you know the formula you can get this answer in a second.

OpenStudy (serenity74):

It's based on which cubic form you are doing. If you are adding: a^3 + b^3 = (a+b) (a^2-ab+b^2) If you are subtracting: a^ - b^3 = (a-b)(a^2+ab+b^2)

OpenStudy (serenity74):

Since you are adding...which formula would you use?

OpenStudy (anonymous):

I would use the first formula.

OpenStudy (serenity74):

*rings bell* Correct answer indeed! :) Let me know if you have anymore questions. Peace.

OpenStudy (anonymous):

Lol ok I will, thanks!

OpenStudy (serenity74):

:)